# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v3

@@ 8 Evaluation of Agentic AI Systems
Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy. We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an AMAS performance and its real-world implications. Below, we discuss these metrics and summarize in Table 7.
Trustworthiness. This criteria evaluates the reliability, safety, and ethical alignment of the AI agent. A trustworthy agent consistently produces correct and unbiased results, adheres to safety constraints, and avoids harmful or unpredictable behavior. Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) [153]. Some approaches combine such factors into an overall trust score. We can define a trustworthiness score $T$ for an Agentic AI system as :
$T=\frac{C+R+I}{S}\,,$ (1)
where $C$ is the agent’s credibility (accuracy and correctness of its outputs), $R$ is reliability (consistency of performance over time), $I$ is the level of user alignment or rapport (analogous to “intimacy” in trust modeling), and $S$ is self-orientation (the degree to which the agent pursues its own goals over the user’s goals). A higher $T$ indicates an agent that is accurate, consistent, user-aligned, and not self-serving, which corresponds to greater trustworthiness. In practice, achieving high trustworthiness means the agent behaves predictably and transparently in accordance with ethical AI principles (such as fairness and accountability).
Explainability. Explainability metrics assess how well the agent’s decisions can be understood and traced by humans [154]. These metrics focus on the clarity and completeness of the rationale the agent provides for its actions. For instance, one can measure the explanation coverage (the percentage of decisions or outputs that come with an adequate explanation) and the fidelity of explanations (how accurately the explanation reflects the true reasoning or model logic).
Consistency of explanations for similar scenarios is another important metric: the agent should explain comparable decisions in a similar way, indicating a stable reasoning process. Quantitatively, methods like OpenXAI [155] provide a suite of metrics to evaluate explanation quality across dimensions such as faithfulness, stability, and fairness of explanations. High explainability builds user trust, as users can follow why the agent made a decision, and it aids debugging by revealing the agent’s internal decision process. In regulated domains (e.g., healthcare or finance), explainability is often essential for compliance and user acceptance.
User-Centered Performance. User-centered metrics capture how effectively the AI agent interacts with and satisfies the end-user’s needs. These criteria emphasize the user’s experience and outcomes [156]. Key metrics include user satisfaction ratings, typically collected via surveys or feedback after interaction, which reflect whether the user’s goals were met and how comfortable they felt with the Agentic AI system behavior.
Additionally, interaction metrics like the number of back-and-forth clarification queries needed (fewer indicates the agent understood the user well) and the coherence or naturalness of the conversational flow contribute to user-centered evaluation. Human-in-the-loop evaluations are often employed here: for example, user studies might rate the agent on criteria such as helpfulness, clarity and naturalness of language, and adherence to user instructions. Ultimately, a user-centered Agentic AI system should align its actions with user intent and preferences.
@@
Table 7: Summary of key evaluation metrics for Agentic AI systems. Each category addresses a different facet of an autonomous agent’s performance and behavior.
@@
Trustworthiness Reliability and safety of the agent’s behavior. Example metrics: task success rate across diverse scenarios; frequency of rule or safety violations (lower is better); calibration of confidence vs. outcomes; fairness and bias indices.
Explainability Transparency and interpretability of the agent’s decisions. Example metrics: percentage of outputs with an accompanying explanation; explanation fidelity to actual model reasoning; consistency of explanations for similar cases; human interpretability ratings of explanations.
User-Centered User satisfaction and alignment with user needs. Example metrics: user satisfaction score (post-interaction survey); rate of fulfilling user-defined goals; number of clarification questions needed (lower is better); naturalness and coherence of dialog from the user’s perspective.
Coordination Effective collaboration in multi-agent or modular systems. Example metrics: multi-agent task completion rate; communication overhead (e.g., messages exchanged); consistency of shared plans or beliefs among agents; synergy score quantifying complementary actions.
Composite Overall performance across metrics. Example metrics: weighted aggregate score encompassing trust, explainability, user-centric, and coordination measures; specialized composite indices (e.g., TUE for tool-use efficacy); cross-domain benchmark results (aggregate success across diverse tasks, as in AgentBench).
Coordination. In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics measure how effectively these parts work together. Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently. Quantitative measures include the team success rate on collaborative tasks (whether the group of agents achieves the overall goal) and communication efficiency metrics (e.g., the number of messages or iterations required among agents to reach a decision, with fewer often indicating more efficient interaction).
One specific example is the Component Synergy Score (CSS), which counts or weights effective interactions between agents, reflecting how well each agent’s actions complement the others (a higher CSS means agents are synergistic rather than working on individual basis). MAS such as ChatDev and MetaGPT provide practical testbeds for these metrics: they orchestrate specialized agents (e.g., different roles in a software engineering pipeline) that must cooperate to complete complex projects.
Evaluations on such frameworks examine whether agents maintain a consistent shared plan, handle inter-agent dependencies smoothly, and recover from misunderstandings. For instance, if one agent generates a plan and another executes it, a coordination metric would assess if the executing agent follows the planner’s intent correctly and whether both agents remain in agreement throughout the process. High coordination scores indicate that the agentic system functions as a cohesive unit, which is crucial for complex tasks beyond the capability of any single agent.
Composite Metrics. Composite metrics aggregate multiple evaluation aspects into a single overall score. These are useful for summarizing agents’ performance holistically, especially when comparing different systems. A composite metric is often a weighted combination of the above categories, for example:
@@
where $M_{T},M_{E},M_{U},M_{C}$ are the normalized scores (on a common scale) for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $w_{T},w_{E},w_{U},w_{C}$ are weights reflecting the relative importance of each aspect for a given application. The choice of weights $w_{i}$ can be domain-specific (for instance, in healthcare applications, trustworthiness and explainability might be weighted more heavily than raw efficiency).
An example of a specialized composite metric in agentic contexts is the Tool Utilization Efficacy (TUE) score, which combines sub-metrics evaluating how correctly and efficiently an agent uses external tools (including proper tool selection and correct parameter usage in tool calls) into one measure. By condensing multiple criteria, composite metrics enable high-level comparison and benchmarking of Agentic AI systems. For instance, AgentBench [157] is a comprehensive benchmark that evaluates agents across a diverse range of tasks and environments (from operating system manipulation to web shopping), effectively providing a composite performance profile of an agent. Such aggregated scores highlight if an agent performs strongly across the board or if it excels in some dimensions while underperforming in others. It is important to interpret composite scores in light of their components: a single number can mask specific weaknesses (e.g., an agent might achieve a high overall score by doing well in task completion and coordination, yet still have poor explainability). Therefore, composite metrics are most informative when accompanied by a breakdown of the agent’s performance per category.
@@ 9 Security and Privacy in Agentic AI Systems
Table 8: Governance Dimensions for Agentic AI Systems
Governance Component Description Example Tools or Methods
Regulatory Compliance Aligning with legal and industry standards for responsible AI [158] NIST AI RMF [99], EU AI Act [159], GDPR [103], HIPAA [104]
Auditability and Logging Capturing traceable records of agent actions and decisions [160] Decision provenance, immutable logs, blockchain-based traceability
Policy Enforcement Enforcing system-level operational, ethical, and security constraints [52, 161] RBAC/ABAC policies [162], formal logic rules, dynamic sandboxing [60, 163, 164]
Human Oversight Human-in-the-loop [140] decision checkpoints and override capabilities Human review agents, interactive dashboards, compliance checkpoints
Risk Monitoring Continuous identification and mitigation of systemic or emergent risks [165] Model risk scanners, out-of-distribution detectors, reinforcement learning audit modules
Explainability Governance Ensuring interpretability standards are met in decision pipelines [166] LIME/SHAP [167, 112, 168], decision trees [169, 170], interpretable surrogate models [171, 172, 173]
Adaptive Governance Updating governance mechanisms based on model evolution and deployment context [133, 135] Governance-as-code [174, 175], auto-updated rule engines [176, 177, 178], learning-based policy adaptation [90, 179]
Incident Response and Recovery Procedures for handling, logging, and recovering from failures or breaches Real-time alerting, kill-switch mechanisms, secure rollback protocols
@@ 9.1 Security Mechanisms
Access control. Access conrol becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [53]. Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [182] and Attribute-Based Access Control (ABAC) [182] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels.
@@ 9.2 Privacy-Preserving Techniques
Data minimization. Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution. For instance, temporary memory buffers used in systems like ChatDev or ReAct-based pipelines are cleared once subgoals are completed, preventing persistent storage of unnecessary user data. Furthermore, anonymization and pseudony [186] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
Secure computation. Techniques including secure multi-party computation (SMPC), homomorphic encryption, and trusted execution environments (TEEs) enable agents to perform computations over encrypted or obfuscated data without compromising privacy. In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage. Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.

### v4

@@ 6 Evaluation of Agentic AI Systems
Agentic AI systems demand evaluation beyond traditional accuracy. We outline five categories: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct facet of performance and real-world impact. Compared to a related work [158], which focuses on proportional impacts across general AI, our work targets Agentic AI and links risks to their causal sources (e.g., prompt injection, memory poisoning). We further introduce an evaluation template with new metrics the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE) to capture multi-agent dynamics. A summary appears in Table 5.
Trustworthiness. This criterion evaluates reliability, safety, and ethical alignment [159]. Core signals include task success rates under distributional shifts (robustness), violation rate of safety/ethical rules (lower is better), and calibration quality (e.g., Expected Calibration Error, ECE). To avoid pathological behavior in scoring, we define a normalized trustworthiness index
$T\;=\;\frac{w_{C}\,C+w_{R}\,R+w_{I}\,I}{\,1+\lambda\,S\,}\,$, (1)
where $C$ (credibility/accuracy), $R$ (reliability over time), and $I$ (user alignment/rapport) are normalized to $[0,1]$, $S\!\geq\!0$ is self-orientation (higher is worse), $w_{C},w_{R},w_{I}\!\geq\!0$ with $w_{C}{+}w_{R}{+}w_{I}\!=\!1$, and $\lambda\!\geq\!0$ controls how strongly self-orientation penalizes trust.
Explainability. Explainability metrics assess how well human evaluators can understand and trace decisions. Useful measures include coverage (fraction of outputs with an explanation), fidelity/faithfulness (agreement with the model’s actual decision basis), stability (consistency across similar cases), and human interpretability ratings. Suites like OpenXAI [160] report faithfulness, stability, and fairness of explanations. High explainability builds user trust and aids debugging; in regulated domains (e.g., healthcare, finance) it is often essential for compliance.
User-Centered Performance. User-centered metrics capture how effectively the agent satisfies end-user needs [161]. Signals include post-interaction satisfaction scores, goal-fulfillment rate, number of clarification turns (lower is better), and perceived coherence/naturalness. Human-in-the-loop (HITL) studies typically rate helpfulness, clarity, and instruction adherence. Ultimately, a user-centered agent aligns actions with user intent and preferences.
@@
Table 5: Summary of key evaluation metrics for agentic AI systems. Each category addresses a distinct facet of performance and behavior.
@@
Trustworthiness Reliability, safety, and alignment. Examples: success rate under shifts; safety/ethics violation rate (lower is better); calibration (ECE/Brier  [^9] https://en.wikipedia.org/wiki/Brier_score); fairness indices; red-team/jailbreak success (lower is better).
Explainability Transparency and traceability. Examples: explanation coverage; faithfulness/fidelity; stability across similar cases; human interpretability scores; explanation time/cost.
User-Centered Experience and outcomes. Examples: satisfaction (Likert/CSAT  [^10] https://www.ibm.com/think/topics/csat-customer-satisfaction-score); goal-completion rate; clarification turns; perceived coherence/naturalness; task time; cognitive load (NASA-TLX).
Coordination Collaboration in multi-agent or modular setups. Examples: team task completion; coordination overhead (messages/tokens, rounds to converge); plan consistency; deadlock/conflict rate (lower is better); synergy score for complementary actions.
Composite Holistic summaries. Examples: weighted aggregate across categories (weights sum to 1); Tool-Use Efficacy (TUE, see definition); cross-domain benchmark aggregates.
Coordination. In multi-agent or modular systems, coordination captures whether components maintain a shared plan, minimize redundant work, and resolve dependencies. Quantify with team success rate, communication efficiency (messages/tokens and rounds to consensus), plan/belief consistency, deadlock/conflict rate, and a Component Synergy Score (CSS) that credits inter-agent actions which enable or improve peers’ outcomes (higher is better).
Composite Metrics.
Composite metrics are often presented as a weighted scalar summary across dimensions:
@@
where $M_{T},M_{E},M_{U},M_{C}\in[0,1]$ are normalized category scores for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $w_{T},w_{E},w_{U},w_{C}\!\geq\!0$ with $\sum w_{\bullet}=1$ encode domain priorities. While this provides a scalar index for ease of interpretation, it risks masking over-performance in one dimension with under-performance in another. To address this, we treat the evaluation more rigorously as a vector of metrics,
$\mathbf{M}\;=\;(M_{T},M_{E},M_{U},M_{C})\,$, (3)
and recommend selection based on $\arg\max$ over $\mathbf{M}$ rather than on any scalarization:
$\arg\max_{i}\;\mathbf{M}_{i}\,.$ (4)
In this view, Eq. 2 can be interpreted as a limiting case, useful for communicating results to non-technical stakeholders, while Eq. 3 and Eq. 4 form the principled basis for evaluation and comparison.
Definition (Tool-Use Efficacy, TUE). In this work, we introduce Tool-Use Efficacy (TUE) as a specialized composite for evaluating how correctly and efficiently an agent uses external tools. Let
[…]
$\mathrm{Exec}$ $\in[0,1]$ $\text{(execution success: no exceptions/timeouts, correct API usage)},$
$\mathrm{Out}$ $\in[0,1]$ $\text{(post-call outcome correctness on tool-gated sub-tasks)},$
$\mathrm{Eff}$ $\in[0,1]$ $\text{(efficiency: normalized inverse cost for latency/tokens/\#calls)}.$
We define two aggregators (both reportable; choose per application):
(a) Linear aggregator (auditable weights).
$\mathrm{TUE}_{\mathrm{lin}}\;=\;\alpha_{\!s}\,\mathrm{Sel}+\alpha_{\!a}\,\mathrm{Arg}+\alpha_{\!e}\,\mathrm{Exec}+\alpha_{\!o}\,\mathrm{Out}+\alpha_{\!f}\,\mathrm{Eff}\,,\quad\alpha_{\!*}\!\geq\!0,\;\sum\alpha_{\!*}=1.$ (5)
(b) Geometric aggregator (“all-of-the-above” strictness).
$\mathrm{TUE}_{\mathrm{geo}}\;=\;\Big{(}\mathrm{Sel}^{\beta_{s}}\,\mathrm{Arg}^{\beta_{a}}\,\mathrm{Exec}^{\beta_{e}}\,\mathrm{Out}^{\beta_{o}}\,\mathrm{Eff}^{\beta_{f}}\Big{)}^{\!1/(\beta_{s}{+}\beta_{a}{+}\beta_{e}{+}\beta_{o}{+}\beta_{f})},\quad\beta_{\!*}\!>\!0.$ (6)
Overhead penalty. If the preference is a separate cost term, then we can define normalized overhead $\mathrm{OH}\in[0,1]$ (increasing in time/tokens/calls) and report the penalized score
$\mathrm{TUE}^{\star}\;=\;\frac{\mathrm{TUE}}{\,1+\lambda\,\mathrm{OH}\,}\,,\quad\lambda\!\geq\!0,$ (7)
where $\mathrm{TUE}$ is either (5) or (6). It is recommended to always report the components (Sel, Arg, Exec, Out, Eff) with confidence intervals to ensure reproducibility.
@@ 7.1 Security Mechanisms
Access control. Access conrol is highly important when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [57]. Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [157] and Attribute-Based Access Control (ABAC) [157] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels.
@@ 7.2 Privacy-Preserving Techniques
Data minimization. Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution. For instance, temporary memory buffers used in systems like ChatDev [30] or ReAct-based pipelines [75] are cleared once subgoals are completed, preventing persistent storage of unnecessary user data. Furthermore, anonymization and pseudony [167] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
Secure computation. Techniques including secure multi-party computation (SMPC) [168], homomorphic encryption [169], and trusted execution environments (TEEs) [^11] Trusted Execution Environment (Wikipedia) enable agents to perform computations over encrypted or obfuscated data without compromising privacy. In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage. Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
@@
Next, we discuss the compliance to AI goverance and policy in Agentic AI systems.
Table 6: Governance controls for agentic AI systems: pillar $\rightarrow$ control $\rightarrow$ operationalization $\rightarrow$ evidence $\rightarrow$ standards.
Pillar Control Operationalization (How) Evidence / Artifacts Standards & Articles
Governance Org. accountability Define owners for models/agents/tools; RACI for changes; approvals for high-impact actions Governance policy; RACI records; change approvals NIST AI RMF (Govern) [171]; ISO/IEC 42001 [2]
Governance Policy-as-code Encode allow/deny rules in orchestrator; pre-exec checks and human sign-off gates Policy repository; policy test cases; gate logs EU AI Act Arts. 11–14 [125]; ISO/IEC 42001 [2]
Risk Mgmt Risk register & AIA Hazard/threat enumeration; likelihood/impact scoring; AI Impact Assessment (AIA) Risk register; AIA reports; mitigation plan ISO/IEC 23894; ISO/IEC 42005 [129]; EU AI Act Art. 9 [125]
Risk Mgmt Post-market monitoring Drift/jailbreak telemetry; incident thresholds; rollback playbooks PMM plan; incident tickets; root-cause analysis EU AI Act Arts. 72–73 [125]; NIST AI RMF (Manage) [171]
Data Gov Data minimization & lineage Task/tenant-scoped memory; TTLs/redaction; dataset lineage and licenses Data inventory; lineage graphs; retention logs EU GDPR Art. 25 [34]; ISO/IEC 5259 series
Data Gov Quality & bias checks Coverage/balance checks; label noise audits; sampling documentation Data quality reports; bias audit results EU AI Act Art. 10 [125]; ISO/IEC TR 24027
Explainability Faithfulness/stability Log rationales; test explanation faithfulness/stability; user-facing summaries XAI eval report; model/data/agent cards NISTIR 8312 [155]; NIST AI RMF (Map/Measure) [171]
Documentation Technical documentation Annex-IV tech docs: purpose, data, performance, oversight, foreseeable misuse Annex-IV dossier; limitations & misuse notes EU AI Act Arts. 11–12; Annex IV [125]
Traceability Provenance & logging Log prompts, plans, tool calls, I/O, timestamps, agent role; sign releases Immutable logs (WORM/tamper-evident); signed artifacts EU AI Act Arts. 12–13 [125]; ISO/IEC 42001 [2]
Human Oversight Human-in-the-loop gates Require human approval for sensitive plans (delete, external write, PII) Approval records; override/rollback logs EU AI Act Art. 14 [125]; NIST AI RMF (Manage) [171]
Robustness Stress/red-team testing Adversarial prompts; perturbation/fault tests; safety regression suites Red-team report; robustness curves; residual risk EU AI Act Art. 15 [125]; ISO/IEC 24029-1 [32]
Security I/O mediation Prompt sanitization; output filtering; sensitive-data (PII/PHI/secrets) detection Filter policies; violation/egress logs OWASP LLM Top-10 [137]; HIPAA §164 [35]
Security Tool access control RBAC/ABAC per agent/role; per-tool API tokens; argument validation; allowlists Access matrices; token scopes; arg-validator tests NIST SP 800-218/218A; NIST AI RMF (Manage)
Security Isolation/sandboxing Containers/VMs; syscall/network/file egress policies; hardware TEEs for sensitive code Sandbox configs; egress policy; attestation logs NIST SP 800-218A; TEEs best-practice notes
Privacy DP budgeting & consent Differential privacy budgets; consent bases; data-subject rights flows DP budget ledger; DSR tickets; consent records GDPR Art. 25 [34]; OECD principles [127]
Supply Chain Provenance & SBOM Track sources and hashes for models/tools/datasets; dependency scanning SBOM; third-party attestations; license checks NIST SSDF (SP 800-218/218A); ISO/IEC 42001
Coordination Role separation & least privilege Distinct agent roles (planner, coder, reviewer); minimal cross-role permissions Role definitions; privilege matrices; audit samples NIST AI RMF (Govern) [171]; ISO/IEC 42001
Coordination Plan consistency checks Shared plan/belief verification; conflict/deadlock detection; CSS metric Plan snapshots; CSS scores; conflict logs Multi-agent eval practice; robustness [32]
Monitoring Anomaly & exfil detection Detect unusual tool chains, volume spikes, strange domains; rate-limit/kill switch Anomaly dashboards; blocked-egress records OWASP LLM Top-10 [137]; NIST AI RMF (Manage)
Transparency User notices & recourse User-facing disclosures, capabilities/limits, and appeal/feedback channels Notices; feedback logs; SLA for appeals OECD principles [127]; EU AI Act transparency
Evaluation Fairness & calibration Group metrics; ECE/Brier; CI reporting; risk–benefit trade-off docs Fairness report; calibration plots; CI tables NIST AI RMF (Measure) [171]; ISO/IEC TR 24027
Records Release/rollback discipline Versioned releases; rollback points; emergency disable/recall Release notes; rollback attestations EU AI Act PMM [125]; ISO/IEC 42001
GPAI GPAI obligations Provide model cards, training-data summary, eval reports, and usage restrictions; disclose known limitations GPAI documentation pack; model card; eval report EU AI Act (GPAI duties) [125]; ISO/IEC 42001
Transparency Synthetic content disclosure Label AI-generated/edited media; attach provenance (e.g., C2PA) and user notices Watermarking/provenance config; disclosure logs EU AI Act Art. 52 [125]; C2PA121212c2pa.org
Privacy Cross-border transfers SCCs/TIAs for third-country transfers; data-residency controls; key management SCC/TIA records; DPA; data-map; key-custody logs GDPR Ch. V [34]
Privacy Data-subject rights ops Verified workflows for access, erasure, rectification, restriction, portability, objection DSR tickets/SLAs; fulfillment proofs; audit samples GDPR Arts. 12–23 [34]
Operations Change mgmt & release gates Risk-based pre-release checklist; dual-control approvals; staged rollouts; rollback criteria Change tickets; gate results; approvals; rollback points ISO/IEC 42001; NIST AI RMF (Manage) [171]
Supply Chain Third-party/vendor risk Vendor questionnaires; DPAs; penetration/attestation reports; license compliance Vendor risk register; DPA file; attestation/SBOM NIST SSDF (SP 800-218/218A); ISO/IEC 42001
Human Factors HCD & accessibility Human-centred design reviews; usability risk analysis; accessibility conformance checks HCD review notes; usability test reports; a11y checklists ISO 9241-210131313ISO 9241-210; OECD principles [127]
Sustainability Energy/CO2 telemetry Track/normalize energy use and CO2; report per release; set budgets Energy/CO2 logs; budget vs. actual; disclosure note Org policy / reporting guidance